OpenStudy (anonymous):

Find the equation of the tangent to the curve x=e^x, y =(t-1)^2, (1,1)

OpenStudy (anonymous):

Slope of tangent is dy/dx Please check x =e^t

OpenStudy (amistre64):

take partials right?

OpenStudy (anonymous):

what does it mean \[x=e^x\]?

OpenStudy (anonymous):

\[x=e ^{t}, y=(t-1)^{2}\]

OpenStudy (anonymous):

we got that the tangent is y=-2x+3 but im not sure what number to evaluate the derivative at

OpenStudy (amistre64):

dx = e^t ; dy = 2(t-1)

OpenStudy (anonymous):

Take partial derivatives dx/dt = e^t dy/dt = 2(t-1) Slope = dy/dx = dy/dt / dx / dt = 2(t-1)/ e^t

OpenStudy (anonymous):

ok i got that far

OpenStudy (anonymous):

but what number do we evaluate that at?

OpenStudy (amistre64):

e^x = 1 (t-1)^2 = 1

OpenStudy (amistre64):

.... typos persist

OpenStudy (anonymous):

(1,1) is given (x,y) is (1,1) or is it t = 1

OpenStudy (anonymous):

its supposed to be t=0 im just not sure where that is coming from

OpenStudy (amistre64):

t = 0

OpenStudy (amistre64):

e^t = 1; when t=0 (t-1)^2 = 1 when t=0

OpenStudy (amistre64):

plug in t=0 into the derivative

OpenStudy (anonymous):

but those parameters wweren't given in the question. are we just supposed to assume when t=0?

OpenStudy (amistre64):

(x=1,y=1) is given

OpenStudy (anonymous):

dy/dx = dy/dt / dx / dt = 2(t-1)/ e^t Find slope at t =0 slope = -2 y = mx +c Plug i n this value And then find c by plugging in (1,1) for (x,y)

OpenStudy (amistre64):

x(t) = 1 when t=0 y(t) = 1 when t=0

OpenStudy (amistre64):

y = -2x +2(1) + 1

OpenStudy (anonymous):

im still unclear where t=0 is coming from

OpenStudy (amistre64):

do you see the point that they give you? that (1,1) in the question?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

(x,y) is the point; which means that (x=1, y=1) x = e^t = 1 y = (t-1)^1 = 1

OpenStudy (amistre64):

^2 ...

OpenStudy (anonymous):

nice

OpenStudy (anonymous):

got it thanks so much!

OpenStudy (amistre64):

youre welcome